A recent graduate took her first job with an investment bank with a signing bonus of $4,000.00. She decides to invest the bonus into a mutual fund that pays 8.00% per year on average. What will be the value of this investment in 20.00 years?
Answer #1

HI

We will use compounded interest formula to find out the value of investment after 20 years

Here Preset Value P = $4,000

Time t = 20 years

r = 8%

Future value F=?

F = P*(1+r)^t

F = 4000*(1+8%)^20 = 4000*1.08^20 = $18,643.83

Thanks